Solve: z(x+2y)p-z(2x+y)q=y^2-x^2


1
Expert's answer
2021-11-18T17:18:01-0500

Solution:

"z(x+2y)p-z(2x+y)q=y^2-x^2;"

The auxiliary equation are:

"\frac{dx}{z(x+2y)}=\frac{dy}{-z(2x+y)}=\frac{dz}{y^2-x^2};"

Considering the first two equations:

"\frac{dx-dy}{3z(x+y)}=0..........(1)"

"dx=dy"

integrating:

"x=y+c_1;"

Considering (1) and the last auxiliary equatoin:

"\frac{dx-dy}{3z(x+y)}=\frac{dz}{x^2+y^2};"

"(x-y)(dx-dy)=3zdz;"

"xdx-xdy-ydx+ydy=3zdz;"

integrating:

"\frac{x^2}{2}-2xy+\frac{y^2}{2}=\frac{3}{2}z^2+c_2;"

"x^2-4xy+y^2=3z^2+c_2;"

Answer:

The solution of the equation is:

"F(x-y,x^2+y^2-3z^2-4xy)=0."